# ExplainReduce: Summarising local explanations via proxies

### v1

@@ Abstract
Most commonly used non-linear machine learning methods are closed-box models, uninterpretable to humans. The field of explainable artificial intelligence (XAI) aims to develop tools to examine the inner workings of these closed boxes. An often-used model-agnostic approach to XAI involves using simple models as local approximations to produce so-called local explanations; examples of this approach include lime, shap, and slisemap. This paper shows how a large set of local explanations can be reduced to a small “proxy set” of simple models, which can act as a generative global explanation. This reduction procedure, ExplainReduce, can be formulated as an optimisation problem and approximated efficiently using greedy heuristics.
@@ 1 Introduction
Explainable artificial intelligence (XAI) aims to elucidate the inner workings of “closed-box” machine learning (ML) models: models that are not readily interpretable to humans. As machine learning has found applications in almost all fields, the need for interpretability has likewise led to the use of XAI in medicine [1], manufacturing [2] and atmospheric chemistry [3], among many other domains. In the past two decades, many different XAI methods have been developed to meet the diverse requirements [4]. These methods produce explanations, i.e., distillations of a closed-box model’s decision patterns. An ideal XAI method should be model-agnostic – applicable to a wide range of model types – and its explanations would be succinct, easily interpretable by the intended user, and stable. Additionally, such explanations would be global, allowing the user to comprehend the entire mechanism of the model. However, producing global explanations is often a challenge. For example, if a model approximates a complex function, describing its behaviour may require describing the complex function itself, thereby defeating the purpose of interpretability.
A common approach to producing model-agnostic explanations is to relax the requirement for explaining the model globally and instead focus on local behaviour [4]. Assuming a degree of smoothness, approximating the closed-box function in a small neighbourhood is often feasible using simple, interpretable functions, such as sparse linear models, decision trees, or decision rules.
We argue that these local explanations are inherently unstable. We initially observed this phenomenon with slisemap [5], where we noted that most items could be accurately approximated with several different local models. In the same vein, take two commonly used local explanation methods, lime [6] and shap [7]. Both methods estimate feature importance locally by averaging predictions from the closed-box model for items sampled near the one being explained. Such explanations can be interpreted as linear approximations for the gradient of the closed-box model. It is well known that the loss landscapes of, e.g., deep neural network models, are not smooth. Therefore, it can be conjectured that the predictions from closed-box can be likewise unstable. Indeed, research has shown that local explanations of neural networks can be manipulated due to the geometry of the closed-box function [8]. Furthermore, numerous variants of lime which aim to increase stability (such as [9, 10]) indicate that instability of lime is a key issue to address. We hypothesise that this instability is a property of all local explanation methods that use simple models to approximate a complex function. Even in the theoretically ideal case, where a complex model is defined by an exact mathematical formula and the local explanations take the form of gradients, there can be points (e.g., sharp peaks) where the gradient is ill-defined, as in Fig. 1. In practice, especially when working with noisy data and averaging over random samples, the ambiguity is exacerbated; one item may have several nearly equally viable explanations. At the same time, some local explanations may accurately approximate many items. This suggests that if we generate a large set of local explanations, we may be able to find a smaller subset thereof, which could effectively replace the full set without sacrificing accuracy. As an added bonus, while the local explanations may be unstable, this subset of explanations may be more stable, as observed in [3].
@@
This paper introduces a procedure, coined ExplainReduce, that can reduce large sets of local explanations to a small subset of so-called proxy models. This small proxy set can be used as a global explanation for the closed-box model, among other possible applications. Fig. 1 provides insight into the process. In the left panel, we show noisy samples (blue dots) from a closed-box function (marked with a solid blue line). The middle panel shows a set of local explanations (as blue and orange dashed lines) produced by the XAI method slisemap [5], with one explanation per data item. The right panel shows a reduction from this large set of local models to two (solid lines), maximising the coverage of items within an error tolerance (shaded area). The procedure offers a trade-off between interpretability, coverage, and fidelity of the local explanation model by finding a minimal subset of local models that can still approximate the closed-box model for most points with reasonable accuracy.
We start with an overview of potential applications of local explanation subsets and introduce related research. We then define the problem of reducing a large set of local models to a proxy set as an optimisation problem. We also outline the ExplainReduce algorithm and our performance metrics. In the results section, we first show how a small proxy set provides a global explanation for the closed-box model with simulated data and a practical example. Second, we find that a proxy set of modest size attains an adherence to the closed-box model on unseen data comparable to, or even better than, the complete set of explanations. We continue to show that we can find a good proxy set even when starting from a limited set of initial explanations. We then demonstrate how greedy approximation algorithms can efficiently solve the problem of finding the proxy set. The code for the procedure and to recreate each of the experiments presented in this paper can be found at https://github.com/edahelsinki/explainreduce.
@@
Figure 1: A simple example of the idea behind ExplainReduce. A closed-box model (left) can have many local explanations (middle). We can reduce the size of the local explanation set to get a global explanation consisting of two simple models (right).
@@ 2 Applications for reduced local explanation sets
Global explanations: Producing succinct global explanations for complex closed-box functions remains a challenge. Given a local explanation method, a naive approach would be simply producing a local explanation for each item in a given training set. However, this leaves the user with $n$ local explanations without guaranteeing global interpretability. In practical settings, many of these local explanations are also likely to be similar and thus redundant. Summarising the large set of local explanations with a reduced set of just a few local models would be much more palatable to the user as a global explanation.
Interpretable replacement for the closed-box model: Assuming that the method used to produce local explanations is faithful, i.e., able to accurately predict the behaviour of the closed-box function locally, the set of all local explanations can replace the closed-box model with little loss to accuracy. However, a large set of (e.g., $500$) local models can hardly be called interpretable. If there is sufficient redundancy in the explanations, selecting a small number of representative models yields a global explanation and provides an interpretable surrogate for the closed-box model.
Exploratory data analysis and XAI method comparison: Studying how the set of local explanations can be reduced can offer interesting insights about the data and XAI method used. If, for example, different subsets of the data are consistently associated with very similar proxy sets, it implies that the closed-box model can be replaced with a reasonable loss of accuracy by a small set of simple functions. Similarly, comparing the reduced proxy sets with individual explanations from different XAI methods could provide new insights into the explanation generation mechanisms in various settings.
Outlier detection: Given a novel observation, we can study how well the local models in the reduced set can predict it compared to similar items in the training dataset. If the novel item is better explained by a different model than the one used for other similar items, or if no explanation in the reduced set accurately captures the relationship, the novel item could be considered an outlier.
@@ 3 Related work
Modern machine learning tools have become indispensable in almost all industry and research fields. As these tools find more and more applications, awareness of their limitations has simultaneously spread. Chief among these is the lack of interpretability inherent to many of the most potent ML methods. Understanding why ML models produce a specific prediction instead of something else poses obstacles in their adoption in, e.g., high-trust applications [11]. Such opaque methods are colloquially called black-box or closed-box methods, in contrast with white- or open-box methods, characterised by their interpretability to humans. Examples of closed-box methods include random forests and deep neural networks, whereas, e.g., statistical models like linear regression models are often considered open-box methods. As a result, the field of XAI has grown substantially in the last decade. This section describes several post-hoc local explanation and model aggregation methods to give the reader context for the ExplainReduce procedure.
@@
XAI methods generally take a dataset and a closed-box function as inputs and produce an explanation that describes the relationship between the inputs and outputs of the closed-box model. One commonly used approach is to produce post-hoc explanations using local (open-box) surrogate models. In this approach, given a closed-box method $f:\mathcal{X}\rightarrow\mathcal{Y},f(\bm{X})=\hat{\bm{y}}$, a local surrogate model $g$ is another function that replicates the behaviour of the closed-box method in some region of the input space. Mathematically, given a loss measure $\ell$ and a point in the input space $\bm{x}_{i}$, we might consider models $g$ such that $\ell(f(\bm{x}^{\prime}_{i}),g(\bm{x}^{\prime}_{i}))\leq\varepsilon\;\forall\bm{x}^{\prime}_{i}\in\{\bm{x}\in\mathcal{X}|D(\bm{x}^{\prime}_{i},\bm{x}_{i})\leq\delta\}$ as local surrogate models.
Many methods have been proposed to produce local surrogate models as explanations. Theoretically, the simplest way to produce a local surrogate model would be to calculate the gradient of the closed-box function. This XAI method is often called vanillagrad in the literature. However, in practice, the gradients of machine learning models can be very noisy, as demonstrated by the effectiveness of adversarial attacks that exploit small perturbations in neural networks [12]. On the other hand, many commonly used machine learning models, such as random forests, do not have well-defined gradients. Hence, more involved approaches are warranted.
smoothgrad [13] attempts to solve the gradient noise problem in vanillagrad by averaging over gradients sampled from the vicinity of the point under explanation. Although the original paper only applies smoothgrad to classification, the method can easily be extended to regression.
Moving away from directly analysing the gradient, lime [6] and shap [7] are perhaps the most widely known examples of practical local explanation generation methods. In this paper, we focus on the kernel-shap variant, which combines linear lime with shap. Both lime and kernel-shap produce explanations in the form of additive linear models ($\hat{y}=\phi^{T}\bm{x}$). Given an item $\bm{x}_{i}$, the methods sample novel items $\bm{x}^{\prime}_{i}\in\mathcal{X}^{\prime}$ in the neighbourhood of the first item, use the closed-box model to predict the labels for the novel items and find the linear model that best fits them [7],
where $\pi_{x}$ represents a distance measure and $\Omega(\phi)$ is a regularisation term. The difference between the methods lies in the choice of distance measure, which defines the notion of neighbourhood for $\bm{x}_{i}$; lime most often uses either $L^{2}$ or cosine distance, whereas kernel-shap utilizes results in game theory [14].
In addition to a procedure for generating local explanations, the authors of lime also propose a method which constructs a global explanation by selecting a set of items whose explanations best capture the global behaviour. They term this procedure submodular pick algorithm. The procedure first generates a local explanation for each item within the dataset. Then, a global feature importance score is calculated as the sum of the square roots of the feature attributions aggregated across all local explanations. If we only pick items according to the presence of the most important features, there is a danger of ending up with many similar explanations. Hence, submodular pick algorithm encourages diversity by framing the problem as a weighted covering problem, balancing feature importance and representativeness such that the user sees a diverse set of explanations.
smoothgrad, lime and shap are based on sampling novel items, which, while simple to implement, introduces a unique set of challenges. First, formulating a reliable data generation process or sampling scheme for all possible datasets is difficult, if not impossible [4, 15]. For example, images generated by adding random noise to the pixel values rarely resemble natural images. Second, randomly generating new items might produce items that cannot occur naturally due to, for example, violating the laws of physics. slisemap [5] and its variant, slipmap [16], produce both a low-dimensional embedding for visualization and a local model for all training items without sampling any new points. slisemap finds both the embedding and local models by optimising a loss function consisting of two parts: an embedding term, where items with similar local explanations attract each other while repelling dissimilar ones and a local loss term for the explanation of each item:
$\min_{g_{i}}\mathcal{L}_{i}=\sum\nolimits_{i=1}^{n}\frac{\exp(-D(\bm{z}_{i},\bm{z}_{j}))}{\sum_{k=1}^{n}\exp(-D(\bm{z}_{k},\bm{z}_{j}))}\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})+\Omega(g_{i})$ (2)
@@
Another large class of explanations which deserves mention is the so-called case-based or example-based XAI methods, which use representative samples in the training set to explain novel ones [17, 18]. One such method is using prototypes, which present the user with the most similar “prototype items” as an explanation, such as showing images of birds with similar plumage as a basis for classification.
A shared property of post-hoc local surrogate models is the lack of uniqueness; for a given item, many local surrogates may exist with similar performance. The phenomenon is documented for slisemap in [5] and implied for other methods based on the results in [8], as well as for many publications aimed at fixing the inherent instability of lime [9, 10, 19]. We argue that the existence of such alternative explanations is an inherent feature of using local surrogate models. Intuitively, we can imagine the $n$-surface of a complex closed-box function and consider local surrogates as planes with dimensionality $n-1$. There are many ways to orient a local surrogate on the curved surface of the closed-box function while retaining reasonable local fidelity. Interestingly, the existence of alternative explanations implies that there may be surrogates which perform well for many items in the data distribution. Therefore, we might be able to reduce a large set of local models to a small set of widely applicable surrogates, providing a global explanation of the model’s behaviour.
The method proposed in the previous paragraph falls under model aggregation. The submodular pick algorithm mentioned when discussing lime is an example of a model aggregation method. Other methods include Global Aggregations of Local Explanations (gale) [20], GLocalX [21], and an integer programming-based approach introduced by Li et al. [22]. gale offers alternative ways to calculate feature importance for the submodular pick algorithm, as the authors argue that the way these importance values are calculated in the original [6] is only applicable to a limited number of scenarios. Furthermore, they show how the choice of the best-performing importance value definition is task-dependent. In GLocalX [21], the authors propose a method to merge rule-based explanations to find a global explanation. In the programming-based approach ([22]), the authors take a similar approach to the one proposed in this paper. They also attempt to find a representative subset of local models, and formulate model aggregation as an optimisation problem with fidelity and coverage constraints. However, their work has some limitations. First, their method relies on the definition of applicability radii for the local models, i.e., radii within which the explanation holds. Second, the framework only functions in classification tasks. Third, to satisfy the optimisation constraints, the framework requires the inclusion of tens of local models into the aggregation, limiting the interpretability of the aggregation as a global model. Finally, they only tested their framework with random forest models and two datasets. Because the model aggregation methods described above cannot be directly applied to an arbitrary set of local explanations, in this paper we opt to measure the performance of ExplainReduce against the full set of local explanations instead.
@@ 4 Methods
In this section, we describe the idea and implementation of ExplainReduce. Assuming that many items in a dataset can have many alternative explanations of similar performance, a proper subset of explanations can accurately model most of the items in the dataset. We refer to this smaller set as a set of “proxy models”. Thus, the method combines aspects of local surrogate explanations with prototype items; instead of representative data items, we use representative local surrogate models to summarise global model performance.
The ExplainReduce procedure works as follows: after training a closed-box model, we generate a large set of local explanations for the closed-box and then find a covering subset of the local models, which acts as a global explanation. In this section, we first define the problem of finding the subset of local models and then move on to cover the reduction methods and algorithms that generate these proxy sets. We also introduce the quality metrics used to evaluate the performance of reduced sets.
@@ 4.1 Problem definition
A dataset $\mathcal{D}=\{(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{n},\bm{y}_{n})\}$ consists $n$ of data items (covariates) $\bm{x}_{i}\in\mathcal{X}$ and labels (responses) $\bm{y}_{i}\in\mathcal{Y}$. We use ${\bm{X}}$ denote a matrix of $n$ rows such that ${\bm{X}}_{i\cdot}=\bm{x}_{i}$. If we have access to a trained supervised learning algorithm $f(\bm{x}_{i})=\hat{\bm{y}}_{i}$, we can instead replace the true labels $\bm{y}_{i}$ with the predictions from the learning algorithm $\hat{\bm{y}}_{i}$. A local explanation for a data item $(\bm{x}_{i},\bm{y}_{i})$ is a simple model $g(\bm{x}_{i})=\tilde{\bm{y}}_{i}$ which locally approximates either the connection between the data items and the labels or the behaviour of the closed-box functions. In the previous section, we gave multiple examples of generating such explanations. Assume we have generated a large set of such local explanations $\bm{G}=\{g_{j}\mid j\in[m]=\{1,\ldots,m\}\}$ and a mapping from data items to models $\Phi:{\mathcal{X}}\mapsto[m]$ using one of these methods. Assume also that these local models can be identified by a set of $p$ real parameters (such as coefficients of linear models), which we will denote with $\bm{B}\in{\mathbb{R}}^{m\times p}$. We define a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbf{R}_{\geq 0}$ and a loss matrix $\bm{L}\in{\mathbb{R}}_{\geq 0}^{m\times n}$, where individual items are defined as $\bm{L}_{ij}=\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})$. A straightforward example of a mapping $\Phi$ would then be chosen from the local models, the one with the lowest loss for each item: $\Phi(\bm{x})=\arg\min\nolimits_{i\in[m]}{\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})}$.
Finally, we assume that the large set $\bm{G}$ contains a reasonable approximation for each item in the dataset ${\cal D}$: for all $i\in[n]$ and for a given $\varepsilon\in{\mathbb{R}}_{>0}$ there exists $j\in[m]$ such that $\ell(g_{j}(\bm{x}_{i}),\bm{y}_{i})\leq\varepsilon$. This can be achieved by, e.g., learning a local explanation for each item in ${\cal D}$.
@@
We are interested in how many items can be explained by a given set of local surrogates to a satisfactory degree. To measure this, we use coverage $C$, defined as the proportion of data items which can be explained sufficiently by at least one local model in a subset ${\bm{S}}\subseteq[m]$. Mathematically, given a loss threshold $\varepsilon\in{\mathbb{R}}_{>0}$, the coverage can be calculated as
$C({\bm{S}},\varepsilon)=(1/n)\left|\{j\in[n]\mid\min\nolimits_{i\in{\bm{S}}}{}\ell(g_{i}(\bm{x}_{j}),\bm{y}_{j})\leq\varepsilon\;\}\right|.$ (3)
A $c$-covering subset of local models ${\bm{S}}\subseteq[m]$ is a set for which $C({\bm{S}},\varepsilon)\geq c$.
@@
Next, we define three computational problems to address the task described earlier. Later, we will introduce algorithms to solve each of the problems.
@@
The first formulation attempts to minimise the number of items for which a satisfactory local explanation is not included in the subset ${\bm{S}}_{c}$.
@@
(maximum coverage) Given $k$ and $\varepsilon\in{\mathbb{R}}_{>0}$, find a subset ${\bm{S}}_{c}$ of cardinality $k$ that maximises coverage, or
${\bm{S}}_{c}=\arg\max\nolimits_{{\bm{S}}\subseteq[m]_{k}}C({\bm{S}},\varepsilon),$ (4)
where we have used $[m]_{k}=\{{\bm{S}}\subseteq[m]\mid\left|{\bm{S}}\right|=k\}$ to denote the subsets of cardinality $k$.
The second definition attempts to capture a subset of explanations that can be used as a proxy model with a small average loss.

### v2

# ExplainReduce: generating global explanations from many local explanations

@@ Abstract
Most commonly used non-linear machine learning methods are closed-box models, uninterpretable to humans. The field of explainable artificial intelligence (XAI) aims to develop tools to examine the inner workings of these closed boxes. An often-used model-agnostic approach to XAI involves using simple models as local approximations to produce so-called local explanations; examples of this approach include lime, shap, and slisemap. This paper shows how a large set of local explanations can be reduced to a small “proxy set” of simple models, which can act as a generative global explanation. This reduction procedure, ExplainReduce, can be formulated as an optimisation problem and approximated efficiently using greedy heuristics. We show that, for many problems, as few as five explanations can faithfully emulate the closed-box model and that our reduction procedure is competitive with other model aggregation methods.
@@ 1 Introduction
Explainable artificial intelligence (XAI) aims to elucidate the inner workings of “closed-box” machine learning (ML) models: models that are not readily interpretable to humans. As machine learning has found applications in almost all fields, the need for interpretability has likewise led to the use of XAI in medicine [band2023medical], manufacturing [peng2022industrial] and atmospheric chemistry [seppalainen2023using], among many other domains. In the past two decades, many different XAI methods have been developed to meet the requirements [guidotti2018survey, dwivedi2023survey] in this diverse set of fields. XAI methods typically produce explanations, i.e., distillations of a closed-box model’s decision patterns. These explanations can be divided into two categories by their applicability in the data space (local vs. global) and by their applicability to different ML models (model-specific vs. model agnostic). Global explanations approximate the general behaviour of a closed-box model everywhere. However, producing such explanations, especially of the model-agnostic variety, is often a challenge: if a model approximates a complex function, describing its behaviour may require describing the complex function itself, thereby defeating the purpose of interpretability. Hence, a common approach to producing model-agnostic explanations is to relax the requirement for explaining the model globally and instead focus on local behaviour [guidotti2018survey, mersha2024survey]. Assuming a degree of smoothness, approximating the closed-box function in a small neighbourhood is often feasible using simple, interpretable functions, such as sparse linear models, decision trees, or decision rules.
This work was motivated by our observation that in practice, the local explanations for similar items will often be similar, provided the underlying closed-box function is smooth in the region. Hence, many local explanations may explain a given item with nearly equal fidelity. We initially observed this phenomenon with slisemap, which embeds the datapoints and their embeddings in a plane based on the similarity of their local explanations. We found that most items could be approximated with almost equal fidelity by using any of the neighbouring local models; see Sect. 4.5 in bjorklund2023slisemap. Conversely, many local models can accurately approximate multiple items. These two observations suggest that many local explanations are redundant. Furthermore, if we generate a large set of local explanations, we may be able to find a smaller subset thereof, which could effectively replace the full set without sacrificing the fidelity of the explanations.
@@
In this paper, we introduce a procedure, coined ExplainReduce, that can reduce large sets of local explanations to a small subset of so-called proxy models. This small proxy set can be used as a global explanation for the closed-box model, among other possible applications. Fig. 1 provides insight into the process. In the left panel, we show noisy samples (blue dots) from a closed-box function (marked with a solid blue line). The middle panel shows a set of local explanations (here linear models, blue and orange dashed lines) produced by the XAI method slisemap [bjorklund2023slisemap], with one explanation per data item. The right panel shows a reduction from this large set of local models to two (solid lines), maximising the coverage of items within an error tolerance (shaded area). The procedure offers a trade-off between interpretability, coverage, and fidelity of the local explanation model by finding a minimal subset of local models that can still approximate the closed-box model for most points with reasonable accuracy, as demonstrated in Section 5. Our approach differs from prior explanation aggregation methods in three ways: the procedure (i) can be applied to both regression and classification tasks, (ii) can be applied to many types of local explanations, and (iii) uses efficient approximate optimisation approaches with an upper bound on worst-case performance (see Section 4).
The contributions of this work are as follows:
  1. 1.
    We introduce ExplainReduce, a model-agnostic procedure for reducing large sets of local explanations into compact proxy sets that serve as global explanations.
  2. 2.
    We formalize explanation reduction as an optimization problem balancing fidelity, coverage, and interpretability, and propose efficient greedy algorithms with theoretical guarantees (Section 4).
  3. 3.
    We empirically demonstrate that hundreds of local post-hoc explanations can be reduced to a representative subset (e.g., five models) while retaining the fidelity of the full explanation set (Section 5).
  4. 4.
    We compare ExplainReduce against existing aggregation methods, showing superior performance and scalability (Section 5.3).
The code for the procedure and to recreate each of the experiments presented in this paper can be found in the supplementary material.
@@
Figure 1: A schematic of the ExplainReduce procedure. A closed-box model (left) can have many local explanations (middle). We can reduce the size of the local explanation set to get a global explanation consisting of two simple models (right).
@@ 2 Applications for reduced local explanation sets
Here, we list some use cases for our proposed algorithm.
Global explanations: Generating one local explanation per data point easily results in hundreds, if not thousands, of explanations, overwhelming users. A reduced proxy set offers a compact, interpretable global view of the closed-box model behaviour.
Interpretable surrogate for the closed-box model: If local explanations faithfully approximate the closed-box model, the full set of these explanations can replace it with minimal accuracy loss. A small proxy set offers a practical, interpretable surrogate without sacrificing accuracy.
Exploratory data analysis and XAI method comparison: Studying reduced explanation sets can reveal redundancy in explanations and highlight clusters of similar models. Comparing proxy sets from different XAI methods offers insights into the stability and consistency of their explanations.
Outlier detection: Given a novel observation, we can assess how well the local models in the reduced set predict it relative to similar items in the training dataset. If the novel item is better explained by a different model than the one used for other similar items, or if no explanation in the reduced set accurately captures the relationship, the novel item could be considered an outlier.
Clustering: By associating each training item with an explanation from a small set, we can partition the data based on the closed-box model behaviour. If the association is achieved, for example, by selecting the explanation with the lowest error, we can perform clustering without a distance measure, which is a desirable property for many types of data for which constructing a distance measure is non-trivial, such as images or molecules.
@@ 3 Related work
Modern machine learning tools have become indispensable in almost all industries and research fields [rashid2024aisurvey]. As these tools find more and more applications, awareness of their limitations has simultaneously spread. Chief among these is the lack of interpretability inherent to many of the most potent ML methods [mersha2024survey, ali2023xai]. Understanding why ML models produce a specific prediction instead of something else poses obstacles in their adoption in, e.g., high-trust applications [dosilovic2018survey]. Such opaque methods are colloquially called black- or closed-box methods, in contrast with white- or open-box methods, characterised by their interpretability to humans. Examples of closed-box methods include random forests and deep neural networks, whereas, e.g., statistical models like linear regression models are often considered open-box methods [molnar:2019:a]. As a result, the field of XAI has grown substantially in the last decade [mersha2024survey].
@@
XAI methods generally take a dataset and a closed-box function as inputs and produce an explanation that describes the relationship between the inputs and outputs of the closed-box model. One commonly used approach is to produce post-hoc explanations using local (open-box) surrogate models [burkhart2021survey]. In this approach, given a closed-box method $f(\bm{X})=\hat{\bm{y}}$, a local surrogate model $g$ is another function that replicates the behaviour of the closed-box method in some region of the input space.
Many methods have been proposed to produce post-hoc local surrogate models as explanations. In this work, we focus mainly on local linear approximations, generated via three general methods. The first category uses gradient information either directly, termed vanillagrad in this article, or by smoothing via sampling the neighbourhood (smoothgrad by smilkov2017smoothgrad). The next category is using the closed-box model as an oracle for items sampled from the neighbourhood of the item under explanation. This category contains lime [ribeiro2016] and shap [lundberg2017unified]. Finally, we consider slisemap [bjorklund2023slisemap] and its variant slipmap [bjorklund2024SLIPMAP], which avoid sampling by simultaneously finding a low-dimensional embedding for the local explanations alongside the explanations themselves. In addition to local linear surrogates, we also consider rule-based explanations, where local explanations take the form of simple decision rules. We use lore [guidotti2019lore] as the exemplary method. For a more detailed description of the XAI methods mentioned above, we refer to Appendix A.
A shared property of post-hoc local surrogate models is the lack of uniqueness; for a given item, many local surrogates may exist with similar performance. The phenomenon is documented for slisemap in [bjorklund2023slisemap] and implied for other methods based on the connection between loss landscapes of neural networks and the instability of explanations [dombrowski2019explanations], as well as for many publications aimed at fixing the inherent instability of lime [shankar2019alime, zafar2019dlime, zhao2021baylime].
Several methods have been proposed to summarise local explanations into a representative set. In their paper describing lime [ribeiro2016], the authors propose an approach called submodular pick, which balances global feature importance scores with model diversity. This limits the method’s applicability to feature-attribution-based explanations. GLocalX [setzu2021glocalx] is a method of merging rule-based explanations to find a global explanation. Consequently, the method is limited in applicability to rule-based explanations for classification models. [li2022optimal] proposes an integer-programming-based approach to finding an optimal set of local explanations. In this approach, the authors aim to maximise the coverage (defined as the number of items in the training set within a radius $r$ of the nearest explanation in the reduced set) while minimising the 0–1 loss of the reduced explanation set. Their approach is also limited to classification only and hinges on a well-defined distance measure in the data space. Furthermore, the complicated IP problem for finding a small explanation set is computationally expensive to solve.
@@
A related paradigm is the so-called case-based or example-based approaches, which summarise data rather than explanations and use representative samples in the training set to explain novel ones [agnar1994casebased, kim2016examples]. One such method is using prototypes, which present the user with the most similar “prototype items” as an explanation, such as showing images of birds with similar plumage as a basis for classification.
ExplainReduce expands the prior work in three key aspects. The method is not limited to either regression or classification problems. It can be applied to a wide set of local XAI methods, provided that the explanation can be used as a surrogate model, i.e., can produce predictions for previously unseen points. Furthermore, we formulate explanation reduction as a set-cover-like optimisation problem with guaranteed worst-case behaviour, and show that our method can find the reduced explanation sets very efficiently using greedy approximations.
@@ 4 Methods
ExplainReduce aims to summarise a large set of local explanations into a small, representative subset – called proxy models – that collectively approximates the behaviour of the original closed-box model. This approach combines ideas from local surrogate explanations and prototype selections: instead of finding representative data items, we select representative explanations.
The procedure consists of three steps: (i) generate a large number of local explanations using any pre-existing XAI method, (ii) formalise the reduction task as an optimisation problem, and (iii) solve the problem using greedy algorithms with theoretical worst-case guarantees.
Below, we define the reduction problem and describe algorithms and evaluation metrics.
@@ 4.1 Problem definition
A dataset $\mathcal{D}=\{(\bm{x}_{1},\bm{y}_{1}),\ldots,(\bm{x}_{n},\bm{y}_{n})\}$ consists $n$ of data items (covariates) $\bm{x}_{i}\in\mathcal{X}$ and labels (responses) $\bm{y}_{i}\in\mathcal{Y}$. We use $\bm{X}$} to denote a matrix of $n$ rows such that ${\bm{X}}_{i\cdot}=\bm{x}_{i}$ when $\cal X$} is a real vector space (as in the examples of this paper). A closed-box model, i.e., a trained supervised learning algorithm $f(\bm{x}_{i})=\hat{\bm{y}}_{i}$, produces predictions from items to labels. It should be noted that much of the analysis in this section applies even if an explicit closed-box model is not provided. In such cases, we treat the data-generating process as the closed-box model ($\bm{\hat{y}}_{i}=\bm{y}_{i}$).
A local explanation for a data item $(\bm{x}_{i},\bm{\hat{y}}_{i})$ is a simple model $g(\bm{x}_{i})=\tilde{\bm{y}}_{i}$ which locally approximates either the connection between the data items and the labels or the behaviour of the closed-box functions. Let $\bm{G}=(g_{1},\dots,g_{m})$ be a set of $m$ such explanations, obtained, e.g., by an XAI model providing local explanations. Next, we define a loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{\geq 0}$ and a loss matrix $\bm{L}\in{\mathbb{R}}_{\geq 0}^{m\times n}$, where individual items are defined as $\bm{L}_{ij}=\ell(g_{i}(\bm{x}_{j}),\bm{\hat{y}}_{j})$. For a given $\varepsilon\in{\mathbb{R}}_{>0}$, we say an explanation $g_{i}$ is faithful for a data item $j$ if $\ell(g_{j}(\bm{x}_{i}),\bm{\hat{y}}_{i})\leq\varepsilon$. Finally, assume that the set of explanations $\bm{G}$ contains at least one faithful approximation for each item in the dataset $\cal D$}. This can be achieved by, e.g., learning a local explanation for each item in $\cal D$}.
@@
We are trying to find an optimal subset of models $\bm{S}\subseteq[m]=\{1,\ldots,m\}$, which we call a proxy set. We want the proxy set $\bm{S}$ to be faithful to the closed-box model. Additionally, we are interested in how many items can be explained by a given set of local surrogates to a satisfactory degree. To measure this, we use coverage $C$, defined as the proportion of data items which can be explained sufficiently by at least one local model in the subset $\bm{S}$. Mathematically, given a loss threshold $\varepsilon\in{\mathbb{R}}_{>0}$, the coverage can be calculated as
$C({\bm{S}},\varepsilon)=\left|\{j\in[n]\mid\min\nolimits_{i\in{\bm{S}}}{}\ell(g_{i}(\bm{x}_{j}),\bm{\hat{y}}_{j})\leq\varepsilon\;\}\right|/n.$ (1)
A $c$-covering subset of local models $\bm{S}$} is a set for which $C({\bm{S}},\varepsilon)\geq c$.
@@
Next, we define three computational problems to address the task described earlier, alongside the algorithms we use to solve these problems.
4.1.1 Maximum coverage.
@@
The first formulation attempts to minimise the number of items for which a satisfactory local explanation is not included in the subset ${\bm{S}}_{c}$. Conceptually, we aim to maximise the probability that for a given item, the proxy set $\bm{S}$ contains at least one local model such that the loss for that item is within the set error tolerance.
@@
(Max Coverage) Given $k$ and $\varepsilon\in{\mathbb{R}}_{>0}$, find a subset ${\bm{S}}_{c}$ of cardinality $k$ that maximises coverage, or
${\bm{S}}_{c}^{k}=\arg\max\nolimits_{{\bm{S}}\subseteq[m]_{k}}C({\bm{S}},\varepsilon),$ (2)
where we have used $m]_{k}=\{{\bm{S}}\subseteq[m]\mid\left|{\bm{S}}\right|=k\}$ to denote the subsets of cardinality $k$.
Problem 1 is a variant of the NP-complete partial set covering problem, sometimes called the max k-cover. The partial set covering problem is an example of submodular maximisation.
We solve Problem 1 using a standard iterative greedy algorithm, which we call reduce (Alg. 1), which takes as an input the cardinality $k$ and the marginal gain metric $\Delta(i\mid\bm{S})$ for which we use the marginal increase of coverage:
$\Delta C(i\mid\bm{S})=C(\bm{S}\cup\{i\},\varepsilon)-C(\bm{S},\varepsilon).$ (3)
Algorithm 1 reduce: Simple greedy optimisation.
Input: $k$: target size; $\Delta(i\mid\bm{S})$: marginal gain function; initialise $\bm{S}\leftarrow\emptyset$.
   Output: $\bm{S}\subseteq[m]$: subset of indices.
procedure reduce( $k,\Delta$)
  while $|\bm{S}|<k$ do
   $i^{*}\leftarrow\operatorname*{arg\,max}_{i\in[m]\setminus\bm{S}}\Delta(i\mid\bm{S})$ ; $\bm{S}\leftarrow\bm{S}\cup\{i^{*}\}$
  end while
return $\bm{S}$
end procedure
The greedy algorithm has a lower bound of achieving coverage at least $1-\left((k-1)/k\right)^{k}$ ($\geq 1-1/e\approx 0.632$) times the optimal solution [nemhauser1978analysis] [^1] Full proof in Appendix H. We refer to this algorithm with Max Coverage.
4.1.2 Minimum loss.
The second definition aims to capture a subset of explanations that can serve as a proxy model with a small average loss. This is equivalent to maximising the faithfulness of the proxy set to the closed-box model.